# Subgradient Convergence Implies Subdifferential Convergence on Weakly Convex Functions: With Uniform Rates Guarantees

### v1

@@ 4 Explicit Uniform Convergence Rates
$\sup_{x\in X\cap\mathcal{O}}\mathbb{H}(\phi(x),\phi_{S}(x))$ (4.1)
@@
Despite their wide applications, tight bounds on the subdifferential difference of these objectives remain largely open in the literature [DD22, Section 5], which motivates our investigations. We will show in Section 5 how our main result in this section, Theorem 4.2, yield tight bounds on the subdifferential differences, and leads to sharp quantitative characterizations on the nonsmooth landscape of stochastic convex-composite formulations in finite samples.
@@ 4.1 Subgradient Selections
Theorem 3.9 shows that, under regularity Assumptions A’ and B’, there is the bound:
$\sup_{x\in X\cap\mathcal{O}}\mathbb{H}(\phi(x),\phi_{S}(x))\leq\sup_{x\in\mathcal{O}}\left\|{G(x)-G_{S}(x)}\right\|$ (4.3)
where $\mathcal{O}$ is an open set. Remarkably, $G$ and $G_{S}$ can be any pair of selections from the subdifferentials (cf. Lemma 3.8 and Theorem 3.9):
@@
This selection process begins with identifying a subgradient for the nonsmooth convex function $h$, which we decompose into a sum $h=h^{sm}+h^{ns}$ of a smooth component $h^{sm}$ and a nonsmooth component $h^{ns}$. To do so, we use $h^{\prime}_{+}$ and $h^{\prime}_{-}$ to denote the right-hand and left-hand derivative of $h$, respectively. We enumerate the set of nondifferentiable points $\{t\in\mathbb{R}:h^{\prime}_{+}(t)\neq h^{\prime}_{-}(t)\}$ as $\{t_{j}\}_{j=1}^{\infty}$, given the fact that these nondifferentiable points are always countable [BV10, Theorem 2.1.2].
For the majority of Section 4, we will work under the assumption that $\sum_{j=1}^{\infty}(h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j}))<\infty$ so that $h^{ns}$ and $h^{sm}$ are well-defined, although this assumption will be unnecessary in our final bound in Theorem 4.2.
@@ 4.2 A Uniform Bound on Subgradient Selections
A key challenge in examining the rate of uniform convergence from $G_{S}$ to $G$ is due to the nonsmooth nature of $G_{S}$ and $G$, whose definition involves indicator functions, as outlined in equation (4.11). To resolve this challenge, our approach leverages the notion of Vapnik–Chervonenkis dimension (VC dimension) [VC71], a concept familiar to experts in statistical learning theory. In short words, VC dimension measures the complexity of sets used in defining these indicator functions (termed as classifiers in statistical learning theory), and this complexity measure governs the uniform generalizability from empirical averages to expectations for these indicator functions.
We state the definition of VC dimension for completeness [Vap13, Section 3.6].
Definition 4.1 (VC Dimension).
@@
The complexity measure ${\rm vc}(\mathcal{F})$ characterizes generalization capabilities of $\mathcal{F}$ and bounds the uniform rate of convergence from $G_{S}$ to $G$. For common techniques of upper bounding VC dimensions of set families, see [vdVW96, Chapter 2]. When the mapping $x\mapsto c(x;\xi)$ are polynomials in $x\in\mathbb{R}^{d}$ for any given $\xi$—a condition met in our applications—we establish a result giving a tight upper bound of the VC dimension, based on results in real algebraic geometry. See Section 6 for details.
Our main results rely on Lipschitz continuity and regularity assumptions concerning the tails of the random vectors $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$. These random vectors appear in our subgradient selection (4.11), and our assumptions about them align with standard treatments in the literature. To state the assumptions, we first recall the concept of a subexponential tail random vector. We define this in terms of Orlicz norms, following [Ver18, Section 3.4.4].
Definition 4.2.
@@
Examples of $\sigma$-sub-exponential random vectors include a normal distribution $Z\sim\mathsf{N}(0,\Sigma)$ where the largest eigenvalue of $\Sigma$ is bounded above by $c\sigma$, or distributions with bounded support, such as $\left\|{Z}\right\|\leq c\sigma$ almost surely, with the numerical constant $c=1/2$ [Ver18, Section 2].
By Markov’s inequality, $\mathbb{P}(|\langle v,Z\rangle|\geq\sigma\cdot u)\leq 2e^{-u}$ for any unit vector $v$ and $u\geq 0$. Furthermore, by Bernstein’s inequality [Ver18, Theorem 2.8.1] and the centering property of sub-exponential random variables [Ver18, Exercise 2.7.10], for i.i.d. $\sigma$-subexponential random variables $Z_{1},Z_{2},\ldots,Z_{m}$, there is the exponential tail bound for the mean $\frac{1}{m}\sum_{i=1}^{m}Z_{i}$ that holds for every $v$ with $\left\|{v}\right\|=1$, and $u\geq 0$:
@@
This assumption is commonly used in empirical process theory to extend high probability bounds from individual points to the uniform control of the supremum of stochastic processes, utilizing the chaining technique. Notably, this assumption naturally holds in a variety of statistical applications. For a reference, see [vdVW96, Chapter 2].
We are now ready to present our main result. The proof of Theorem 4.1, which relies on tools from statistical learning theory, is relatively standard, and thus deferred to Section A.
Theorem 4.1.
@@
Recall $\mathbb{P}^{*}$ denotes the outer probability (Section 2.4), also used to address measurability issues with supremums in empirical process theory [vdVW96]. This is required because the supremum in equation (4.14) may not be measurable under the probability space.
Theorem 4.1 establishes a uniform closeness between the selected subgradient $G_{S}$ and $G$ over the Euclidean ball $\mathbb{B}(x_{0};r)$. In the final bound (4.14), $\zeta$ captures how the nondifferentiability of $h$ could potentially deteriorate the convergence rate. Crucially, the error term $\Delta$ in equation (4.16) demonstrates how the dimensionality $d$ and the complexity measure ${\rm vc}(\mathcal{F})$ of the set family $\mathcal{F}$ influence the uniform convergence rate.
The presence of $\max\{\Delta,\Delta^{2}\}$ in the final bound instead of just $\Delta$ stems from the sub-exponential assumption (cf. $\min\{u,u^{2}\}$ in the Bernstein inequality (4.13)). When $\Delta\leq 1$, which is typical for large sample size $m$, this term simplifies to $\Delta$, which is $\sqrt{(d+{\rm vc}(\mathcal{F}))/m}$ up to logarithmic factors. Our convergence rates align well with the standard uniform convergence rate $\sqrt{d/m}$ for averages of smooth functions and $\sqrt{{\rm vc}(\mathcal{F})/m}$ for indicator functions in the statistics literature [Wai19]. This suggests our bounds are in general tight up to logarithmic factors in $m$.
Finally, the $\log m$ term arises from the subexponential tail assumption and can be removed if a norm boundedness assumption replaces it in Assumption C.2, or by using more advanced arguments involving additional probabilistic control of the envelope function of $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$ [vdVW96, Chapter 2]. For conciseness and clarity, this paper does not pursue such removal.
@@ 4.3 Implication: A Uniform Bound on Subdifferentials
We integrate Theorem 3.9 and Theorem 4.1 to document our final result, Theorem 4.2, on uniform convergence of subdifferentials. The only work is to streamline the assumptions, recognizing that Assumptions C.1 and C.2 essentially imply that Assumption B’ holds for $\mathcal{O}=\mathbb{B}(x_{0};r)$.
Theorem 4.2.
Assume Assumptions C.1C.2, and $f(x)=\mathbb{E}_{\xi\sim P}[h(c(x;\xi))]<\infty$ for $x\in\mathbb{B}(x_{0};r)$.
@@
In the above, $\zeta,\Delta$ follows the same definition in the statement of Theorem 4.1.
@@
Given Theorem 3.9 and Theorem 4.1, and with Assumption A’ assumed for $\mathcal{O}=\mathbb{B}(x_{0};r)$ in the statement of Theorem 4.2, the remaining task is to show that Assumption B’ is met for $\mathcal{O}=\mathbb{B}(x_{0};r)$. This can be achieved by first recognizing that for $y\in\mathbb{B}(x_{0};r)$:
@@
Assumptions C.1C.2 imply that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)}\right\|$ and $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{(h^{sm})^{\prime}(c(y;\xi))\nabla c(y;\xi)}\right\|$ are integrable over $\xi\sim\mathbb{P}$ due to the standard chaining argument [Ver18, Section 8]. This implies that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)\partial h(c(y;\xi))}\right\|$ is also integrable over $\xi$ since $\sum_{j=1}^{\infty}(h_{+}^{\prime}(t_{j})-h_{-}^{\prime}(t_{j}))<\infty$, thus confirming the first condition of Assumption B’ holds for $\mathcal{O}=\mathbb{B}(x_{0};r)$. The second condition of Assumption B’ follows trivially from Assumption C.2. ∎
Remark 4.3.
Theorem 4.2 establishes a rate for uniform convergence of subdifferential mappings for stochastic convex-composite objectives at $\sqrt{\max\{d,{\rm vc}(\mathcal{F})\}/m}$, modulo logarithmic factors. Our results are different from approaches that depend on the continuity of both empirical and population subdifferentials under the Hausdorff metric, or require a nonatomic probability distribution $\mathbb{P}$ to achieve a $1/\sqrt{m}$ rate [SX07, RX11] (see Section 1.1 for a more detailed literature review). Indeed, Theorem 4.2 achieves this tight rate without requiring continuity in the subdifferential mappings of the objective or specific conditions on the data distribution $\mathbb{P}$.
In the literature on uniform convergence rates of subdifferential mappings for stochastic weakly-convex objectives, methods using Moreau envelope smoothing [Mor65] and Attouch’s epigraphical convergence theorem [Att77] achieve a convergence rate of $4]{1/m}$ suboptimal in terms of the sample size $m$ [DDP20, DD22]. This $4]{1/m}$ convergence is measured using graphical distance, a notion which is topologically stronger than the uniform convergence criterion we consider in Theorem 4.2 (see Remark of Theorem 3.1). Our work complements these studies by showing a convergence rate’s dependence on sample size $m$ to be $\sqrt{1/m}$ (modulo logarithmic factors) under a topologically weaker notion of convergence.
@@ 5 Applications
In this section, we return to some of the examples previously discussed in Section 3.3 to demonstrate the application of our main Theorem 4.2. We will concretely evaluate the bounds in Theorem 4.2 for some of these objectives, compare these to existing results in the literature, and use them to derive new findings.
@@ 5.1 Illustration I: Phase Retrieval
We start with the (real-valued) robust phase retrieval objective (Example 3.5):
@@
This objective has been studied in the literature [EM14, DR19, DDP20], yet tight bounds on the gap between the empirical and population subdifferential maps have yet to be established.
@@
We need to specify conditions on the distribution of measurements $(a,b)\sim\mathbb{P}$ that allow us to apply Theorem 4.2. We recall the concept of a subgaussian tail random vector [Ver18, Section 2.5].
Definition 5.1.
@@
Examples of a $\sigma$-subgaussian distribution include normal distributions $Z\sim\mathsf{N}(0,(c\sigma)^{2}I)$, and uniform distribution on the hypercube $Z\sim{\rm unif}\{\pm(c\sigma)\}^{d}$ where $c>0$ can be any constant obeying $c\leq 1/4$. Applying Theorem 4.2, we derive an upper bound on the subdifferential gap for the robust phase retrieval. Remarkably, our result, which achieves an almost optimal $m^{-1/2}$ rate (up to logarithmic factors), holds regardless of whether the random variables $a$ and $b$ are discrete or continuous. This contrasts with existing literature, which either requires a continuity assumption or attains a slower $m^{-1/4}$ rate (see Remark of Theorem 4.2).
Theorem 5.1.
@@
We apply Theorem 4.2. We first check the assumptions. For the nonsmooth $h(z)=|z|$, it is decomposed as $h=h^{ns}+h^{sm}$ where $h^{sm}(z)=-z$ and $h^{ns}(z)=2(z)_{+}$.
@@
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Notably, for every vector $v$, $\langle a,v\rangle$ is $\left\|{v}\right\|$-subgaussian. Given the property that the product of two subgaussian random variables is subexponential [Ver18, Lemma 2.7.7], we derive that for some universal constant $C>0$, each of
@@
We then compute the bound in Theorem 4.2. Namely, we need to compute $\zeta$ and ${\rm vc}(\mathcal{F})$.
@@
We now bound its VC dimension using Theorem 6.1, whose proof is based on a beautiful theorem in real algebraic geometry [Mil64, Tho65, Mat13]. Note $x\mapsto\langle a,x\rangle^{2}-b$ is a degree $2$ polynomial in $x\in\mathbb{R}^{d}$ for every $a,b$. Using Theorem 6.1, we obtain that for some universal constant $C>0$:
@@
Given $r>0$, we apply Theorem 6.1 to the open ball $\mathbb{B}(0;2r)$, and we obtain that for some universal constant $C>0$:
@@
Theorem 5.1 then follows by recognizing the set inclusion: ${\rm cl}(\mathbb{B}(0;r))\subseteq\mathbb{B}(0;2r)$. ∎
In the literature on optimization, statistics and machine learning, establishing a uniform upper bound on the convergence of subdifferentials is important because it facilitates our understanding of the landscape of the nonsmooth, nonconvex finite sample objective $\phi_{S}$, e.g., the location of stationary points, growth conditions, etc., through the lens of the population objective $\phi$. We shall give further applications in Section 5.1.1 how Theorem 5.1 contributes to existing results to refine our understanding of the landscape of the nonsmooth phase retrieval objective $\Phi_{S}$ in finite samples.

### v2

@@ 4 Explicit Uniform Convergence Rates
$\sup_{x\in X\cap\mathcal{O}}\mathbb{H}(\partial\phi(x),\partial\phi_{S}(x))$ (4.1)
@@
Despite their wide applications, tight bounds on the subdifferential difference of these objectives remain largely open in the literature [16, Section 5], which motivates our investigations. We will show in Section 5 how our main result in this section, Theorem 5, yields tight bounds on the subdifferential differences, and leads to sharp quantitative characterizations on the nonsmooth landscape of stochastic convex-composite formulations in finite samples.
@@ 4.1 Subgradient Selections
Theorem 3 shows that, under regularity Assumptions A’ and B’, there is the bound:
$\sup_{x\in X\cap\mathcal{O}}\mathbb{H}(\partial\phi(x),\partial\phi_{S}(x))\leq\sup_{x\in\mathcal{O}}\left\|{G(x)-G_{S}(x)}\right\|$ (4.3)
where $\mathcal{O}$ is an open set. Remarkably, $G$ and $G_{S}$ can be any pair of selections from the subdifferentials (cf. Lemma 2 and Theorem 3):
@@
This selection process begins with identifying a subgradient for the nonsmooth convex function $h$, which we decompose into a sum $h=h^{sm}+h^{ns}$ of a smooth component $h^{sm}$ and a nonsmooth component $h^{ns}$. To do so, we use $h^{\prime}_{+}$ and $h^{\prime}_{-}$ to denote the right-hand and left-hand derivative of $h$, respectively. We enumerate the set of nondifferentiable points $\{t\in\mathbb{R}:h^{\prime}_{+}(t)\neq h^{\prime}_{-}(t)\}$ as $\{t_{j}\}_{j=1}^{\infty}$, given the fact that these nondifferentiable points are always countable [42, Theorem 2.1.2].
For the majority of Section 4, we will work under the assumption that $\sum_{j=1}^{\infty}(h^{\prime}_{+}(t_{j})-h^{\prime}_{-}(t_{j}))<\infty$ so that $h^{ns}$ and $h^{sm}$ are well-defined, although this assumption will be unnecessary in our final bound in Theorem 5.
@@ 4.2 A Uniform Bound on Subgradient Selections
A key challenge in examining the rate of uniform convergence from $G_{S}$ to $G$ is due to the nonsmooth nature of $G_{S}$ and $G$, whose definition involves indicator functions, as outlined in equation (4.11). To resolve this challenge, our approach leverages the notion of Vapnik–Chervonenkis dimension (VC dimension) [43], a concept familiar to experts in statistical learning theory. In short words, VC dimension measures the complexity of sets used in defining these indicator functions (termed as classifiers in statistical learning theory), and this complexity measure governs the uniform generalizability from empirical averages to expectations for these indicator functions.
We state the definition of VC dimension for completeness [22, Section 3.6].
Definition 1 (VC Dimension).
@@
The complexity measure ${\rm vc}(\mathcal{F})$ characterizes generalization capabilities of $\mathcal{F}$ and bounds the uniform rate of convergence from $G_{S}$ to $G$. For common techniques of upper bounding VC dimensions of set families, see [31, Chapter 2]. When the mapping $x\mapsto c(x;\xi)$ are polynomials in $x\in\mathbb{R}^{d}$ for any given $\xi$—a condition met in our applications—we establish a result giving a tight upper bound of the VC dimension, based on results in real algebraic geometry. See Section 6 for details.
Our main results rely on Lipschitz continuity and regularity assumptions concerning the tails of the random vectors $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$. These random vectors appear in our subgradient selection (4.11), and our assumptions about them align with standard treatments in the literature. To state the assumptions, we first recall the concept of a subexponential tail random vector. We define this in terms of Orlicz norms, following [44, Section 3.4.4].
Definition 2.
@@
Examples of $\sigma$-sub-exponential random vectors include a normal distribution $Z\sim\mathsf{N}(0,\Sigma)$ where the largest eigenvalue of $\Sigma$ is bounded above by $c\sigma$, or distributions with bounded support, such as $\left\|{Z}\right\|\leq c\sigma$ almost surely, with the numerical constant $c=1/2$ [44, Section 2].
By Markov’s inequality, $\mathbb{P}(|\langle v,Z\rangle|\geq\sigma\cdot u)\leq 2e^{-u}$ for any unit vector $v$ and $u\geq 0$. Furthermore, by Bernstein’s inequality [44, Theorem 2.8.1] and the centering property of sub-exponential random variables [44, Exercise 2.7.10], for i.i.d. $\sigma$-subexponential random variables $Z_{1},Z_{2},\ldots,Z_{m}$, there is the exponential tail bound for the mean $\frac{1}{m}\sum_{i=1}^{m}Z_{i}$ that holds for every $v$ with $\left\|{v}\right\|=1$, and $u\geq 0$:
@@
This assumption is commonly used in empirical process theory to extend high probability bounds from individual points to the uniform control of the supremum of stochastic processes, utilizing the chaining technique. Notably, this assumption naturally holds in a variety of statistical applications. For a reference, see [31, Chapter 2].
We are now ready to present our main result. The proof of Theorem 4 employs empirical process theory tools, namely, the calculus rules for outer probability measure [31, Chapter 1.2-5], the chaining argument [45], and the Sauer-Shelah Lemma—a fundamental combinatorial principle using the VC dimension to control function class complexities [46, 47]. The proof is relatively standard for experts in statistical learning theory and is detailed in Section A.
Theorem 4.
@@
Recall $\mathbb{P}^{*}$ denotes the outer probability (Section 2.4), also used to address measurability issues with supremums in empirical process theory [31]. This is technically required because the supremum in equation (4.14) may not be measurable under the probability space.
Theorem 4 establishes a uniform closeness between the selected subgradient $G_{S}$ and $G$ over the Euclidean ball $\mathbb{B}(x_{0};r)$. In the final bound (4.14), $\zeta$ captures how the nondifferentiability of $h$ could potentially deteriorate the convergence rate. Crucially, the error term $\Delta$ in equation (4.16) demonstrates how the dimensionality $d$ and the complexity measure ${\rm vc}(\mathcal{F})$ of the set family $\mathcal{F}$ influence the uniform convergence rate.
The presence of $\max\{\Delta,\Delta^{2}\}$ in the final bound instead of just $\Delta$ stems from the sub-exponential assumption (cf. $\min\{u,u^{2}\}$ in the Bernstein inequality (4.13)). When $\Delta\leq 1$, which is typical for large sample size $m$, this term simplifies to $\Delta$, which is $\sqrt{(d+{\rm vc}(\mathcal{F}))/m}$ up to logarithmic factors. Our convergence rates align well with the standard uniform convergence rate $\sqrt{d/m}$ for averages of smooth functions and $\sqrt{{\rm vc}(\mathcal{F})/m}$ for indicator functions in the statistics literature [48]. This suggests our bounds are in general tight up to logarithmic factors in $m$.
Finally, the $\log m$ term arises from the subexponential tail assumption and can be removed if a norm boundedness assumption replaces it in Assumption C.2, or by using more advanced arguments involving additional probabilistic control of the envelope function of $\nabla c(x,\xi)$ and $(h^{sm})^{\prime}(c(x,\xi))\cdot\nabla c(x,\xi)$ [31, Chapter 2]. For conciseness and clarity, this paper does not pursue such removal.
@@ 4.3 Implication: A Uniform Bound on Subdifferentials
We integrate Theorem 3 and Theorem 4 to document our ultimate result, Theorem 5, on uniform convergence of subdifferentials. At this point, the only work is to collect and streamline all the assumptions in Theorem 3 and 4. To fulfill Assumption B’ required in Theorem 3, we identify an additional condition beyond Assumptions C.1C.2. This Lipschitz-type condition is very mild, and holds in many applications in stochastic programming [2, Chapter 5] and statistics [49, Chapter 5].
Assumption C.3.
There is a measurable function $L:\Xi\to\mathbb{R}$ with $\mathbb{E}_{\xi\sim\mathbb{P}}[L(\xi)]<\infty$ such that
$\begin{split}\left\|{\nabla c(x_{1},\xi)-\nabla c(x_{2},\xi)}\right\|&\leq L(\xi)\left\|{x_{1}-x_{2}}\right\|\\ \left\|{e(x_{1},\xi)-e(x_{2},\xi)}\right\|&\leq L(\xi)\left\|{x_{1}-x_{2}}\right\|\end{split}$
hold for every $x_{1},x_{2}\in\mathbb{B}(x_{0};r)$, where $e(x,\xi)=(\psi^{sm})^{\prime}(c(x,\xi))\nabla c(x,\xi)$.
Theorem 5 documents the final result on uniform convergence of subdifferentials, recognizing that Assumptions C.1C.3 imply that Assumption B’ holds for $\mathcal{O}=\mathbb{B}(x_{0};r)$.
Theorem 5.
Assume Assumptions C.1C.3, and $f(x)=\mathbb{E}_{\xi\sim P}[h(c(x;\xi))]<\infty$ for $x\in\mathbb{B}(x_{0};r)$.
@@
In the above, $\zeta,\Delta$ follows the same definition in the statement of Theorem 4.
@@
Given Theorem 3 and Theorem 4, and with Assumption A’ assumed for $\mathcal{O}=\mathbb{B}(x_{0};r)$ in the statement of Theorem 5, the remaining task is to show that Assumption B’ is met for $\mathcal{O}=\mathbb{B}(x_{0};r)$. This can be achieved by first recognizing that for $y\in\mathbb{B}(x_{0};r)$:
@@
Assumptions C.1C.2 imply that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)}\right\|$ and $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{(h^{sm})^{\prime}(c(y;\xi))\nabla c(y;\xi)}\right\|$ are integrable over $\xi\sim\mathbb{P}$ due to the standard chaining argument [44, Section 8]. This implies that $\sup_{y\in\mathbb{B}(x_{0};r)}\left\|{\nabla c(y;\xi)\partial h(c(y;\xi))}\right\|$ is also integrable over $\xi$ since $\sum_{j=1}^{\infty}(h_{+}^{\prime}(t_{j})-h_{-}^{\prime}(t_{j}))<\infty$, thus confirming the first condition of Assumption B’ holds for $\mathcal{O}=\mathbb{B}(x_{0};r)$. The second condition of Assumption B’ follows trivially from Assumption C.3. ∎
Remark C.
Theorem 5 establishes a rate for uniform convergence of subdifferential mappings for stochastic convex-composite objectives at $\sqrt{\max\{d,{\rm vc}(\mathcal{F})\}/m}$, modulo logarithmic factors. Our results are different from approaches that depend on the continuity of both empirical and population subdifferentials under the Hausdorff metric, or require a nonatomic probability distribution $\mathbb{P}$ to achieve a $1/\sqrt{m}$ rate [1, 21] (see Section 1.1 for a more detailed literature review). Indeed, Theorem 5 achieves this tight rate without requiring continuity in the subdifferential mappings of the objective or specific conditions on the data distribution $\mathbb{P}$.
In the literature on uniform convergence rates of subdifferential mappings for stochastic weakly-convex objectives, methods using Moreau envelope smoothing [23] and Attouch’s epigraphical convergence theorem [15] achieve a convergence rate of $4]{1/m}$ suboptimal in terms of the sample size $m$ [14, 16]. This $4]{1/m}$ convergence is measured using graphical distance, a notion which is topologically stronger than the uniform convergence criterion we consider in Theorem 5 (see Remark of Theorem 1). Our work complements these studies by showing a convergence rate’s dependence on sample size $m$ to be $\sqrt{1/m}$ (modulo logarithmic factors) under a topologically weaker notion of convergence.
@@ 5 Applications
In this section, we return to some of the examples previously discussed in Section 3.3 to demonstrate the application of our main Theorem 5. We will concretely evaluate the bounds in Theorem 5 for some of these objectives, compare these to existing results in the literature, and use them to derive new findings.
@@ 5.1 Illustration I: Phase Retrieval
We start with the (real-valued) robust phase retrieval objective (Example 1):
@@
This objective has been studied in the literature [37, 8, 14], yet tight bounds on the gap between the empirical and population subdifferential maps have yet to be established.
@@
We need to specify conditions on the distribution of measurements $(a,b)\sim\mathbb{P}$ that allow us to apply Theorem 5. We recall the concept of a subgaussian tail random vector [44, Section 2.5].
Definition 3.
@@
Examples of a $\sigma$-subgaussian distribution include normal distributions $Z\sim\mathsf{N}(0,(c\sigma)^{2}I)$, and uniform distribution on the hypercube $Z\sim{\rm unif}\{\pm(c\sigma)\}^{d}$ where $c>0$ can be any constant obeying $c\leq 1/4$. Applying Theorem 5, we derive an upper bound on the subdifferential gap for the robust phase retrieval. Remarkably, our result, which achieves an almost optimal $m^{-1/2}$ rate (up to logarithmic factors), holds regardless of whether the random variables $a$ and $b$ are discrete or continuous. This contrasts with existing literature, which either requires a continuity assumption or attains a slower $m^{-1/4}$ rate (see Remark of Theorem 5).
Theorem 6.
@@
We apply Theorem 5. We first check the assumptions. For the nonsmooth $h(z)=|z|$, it is decomposed as $h=h^{ns}+h^{sm}$ where $h^{sm}(z)=-z$ and $h^{ns}(z)=2(z)_{+}$.
@@
    where $e(x,\xi)=(h^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Notably, for every vector $v$, $\langle a,v\rangle$ is $\left\|{v}\right\|$-subgaussian. Given the property that the product of two subgaussian random variables is subexponential [44, Lemma 2.7.7], we derive that for some universal constant $C>0$, each of
@@
  • (Assumption C.3). It is easy to verify that
    $\begin{split}\left\|{\nabla c(x_{1},\xi)-\nabla c(x_{2},\xi)}\right\|&=2\left\|{\langle a,x_{1}-x_{2}\rangle a}\right\|\leq 2\left\|{a}\right\|^{2}\left\|{x_{1}-x_{2}}\right\|\\ \left\|{e(x_{1},\xi)-e(x_{2},\xi)}\right\|&=2\left\|{\langle a,x_{1}-x_{2}\rangle a}\right\|\leq 2\left\|{a}\right\|^{2}\left\|{x_{1}-x_{2}}\right\|\end{split},$
    where $e(x,\xi)=(\psi^{sm})^{\prime}(c(x;\xi))\nabla c(x;\xi)$. Since $a$ is $\sigma$-subgaussian, $\mathbb{E}[L(\xi)]=2\mathbb{E}[\left\|{a}\right\|^{2}]\leq C\sigma^{2}d$ for some universal constant $C>0$. Thus Assumption C.3 is satisfied with $L(\xi)=2\left\|{a}\right\|^{2}$.
@@
We then compute the bound in Theorem 5. Namely, we need to compute $\zeta$ and ${\rm vc}(\mathcal{F})$.
@@
We now bound its VC dimension using Theorem 8, whose proof is based on a beautiful theorem in real algebraic geometry [50, 51, 52]. Note $x\mapsto\langle a,x\rangle^{2}-b$ is a degree $2$ polynomial in $x\in\mathbb{R}^{d}$ for every $a,b$. Using Theorem 8, we obtain that for some universal constant $C>0$:
@@
Given $r>0$, we apply Theorem 8 to the open ball $\mathbb{B}(0;2r)$, and we obtain that for some universal constant $C>0$:
@@
Theorem 6 then follows by recognizing the set inclusion: ${\rm cl}(\mathbb{B}(0;r))\subseteq\mathbb{B}(0;2r)$. ∎
In the literature on optimization, statistics and machine learning, establishing a uniform upper bound on the convergence of subdifferentials is important because it facilitates our understanding of the landscape of the nonsmooth, nonconvex finite sample objective $\phi_{S}$, e.g., the location of stationary points, growth conditions, etc., through the lens of the population objective $\phi$. We shall give further applications in Section 5.1.1 how Theorem 6 contributes to existing results to refine our understanding of the landscape of the nonsmooth phase retrieval objective $\Phi_{S}$ in finite samples.